OpenStudy (anonymous):

Solve for x: 2/5 x + 2 = 10 x = 20 x = 24 x = 4 x = 8

OpenStudy (rmrjr22):

easy peasy

OpenStudy (rmrjr22):

\[2/5x+2=10......2/5x =10-2\]

OpenStudy (rmrjr22):

from there get rid of the denominator.. u know how?

OpenStudy (anonymous):

2/5x=8 to get x by itself flip the fraction and multiply and then simplify the final fraction

OpenStudy (rmrjr22):

2x=8(5)

OpenStudy (rmrjr22):

x= (8(5))/2... solve this

OpenStudy (anonymous):

20

OpenStudy (rmrjr22):

bam!